# Solving Harder Reinforcement Learning Problems

### v1

@@ 45 N = len(g)
(c) All $k$ sized subsets with $k-1$ sized subsets baseline: $s^{(\mathrm{loo}-1)}$
@@
Figure 7: The effect of the LOO baseline on the effective rewards derived from $n=32$ raw rewards $g(x_{i})$ sampled uniformly from $[-1/2,+1/2]$. The un-baselined effective rewards (a) from (19) include a vertical offset that grows with $k$ despite being a function of raw rewards (horizontal axis) that are centered around zero. The baselined effective rewards (b) and (c) from (29) and (33) respectively are more centered, and give rise to reduced gradient estimator variance. To construct the figure we grouped reward values into regularly spaced bins and averaged the transformed reward for each bin to construct the curves.
Figure 8: The variance of different estimators of the gradient of $\mathrm{max}_{g}\mathrm{@k}$ with $k=4$ for the one-dimensional problem depicted in Figure 1 at the location $x=1$. Each data-point is the sample variance of 10,000 independent unbiased gradient estimates (lower is better). The horizontal axis denotes the number of samples $n$ used to construct each of the 10,000 estimates. We compare the following methods:
all subsets baselined: $s^{(\mathrm{loo})}$ — our novel estimator of Equation 29 that analytically sums over all subsets of size $k$ of the $n$ samples with our unbiased baseline method that subtracts for each element $i$ the mean of the estimator over all subsets of size $k$ that do not include $i$.
naive partitioned baselined — a naive transformation that sets all $k$ transformed rewards in a subset of $k$ samples equal to the largest raw reward in that subset. To extend this method to $n>k$ we partition the $n$ samples (for integer multiples of $k$) into disjoint subsets of size $k$ and average the estimated gradient obtained from each. Furthermore, as a simple variance reduction method, for each such set of $k$ samples we subtract the mean of the transformed rewards from the other sets of $k$ samples (thereby averaging over $(n-k)$ samples and subtracting the result from the $k$ samples and repeating $n/k$ times in a leave-one-out fashion over the subsets of size $k$). If we were to randomly sample an increasing number of partitions of the samples and average over all of them, then intuitively the resulting estimator would approach the variance of $s^{(\mathrm{loo})}$, but this would be expensive and indeed the limiting case of considering all partitions is intractable for general $n$ and $k$. Our estimators have the key property of summing over all such partitions while nonetheless being efficient to compute.
loo minus one all subsets: $s^{(\mathrm{loo}-1)}$ — our novel estimator of Equation 33 that analytically sums over all subsets of size $k$ of the $n$ samples and uses all appropriate subsets of size $k-1$ to form the variance-reducing baseline that retains unbiasedness, thereby non-trivially generalizing [32] to all $n>k$ with strong variance reduction.

### v2

@@ Appendix C Additional Figures
(c) All $k$ sized subsets with k1k-1 sized subsets baseline: $s^{(\mathrm{loo}-1)}$
@@
Figure 5: The effect of the LOO baseline on the effective rewards derived from $n=32$ raw rewards $g(x_{i})$ sampled uniformly from [1/2,+1/2][-1/2,+1/2]. The non baselined effective rewards (a) from (19) include a vertical offset that grows with $k$ despite being a function of raw rewards (horizontal axis) that are centered around zero. The baselined effective rewards (b) and (c) from (29) and (33) respectively are more centered, and give rise to reduced gradient estimator variance. To construct the figure we grouped reward values into regularly spaced bins and averaged the transformed reward for each bin to construct the curves. Note: because our transformations are from $\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ it is not possible to directly inspect a one-dimensional transformation.
(a) Cumulative solve rate ($N=12000$ problems)
(b) Effect of changing $k$ on entropy during training
Figure 6: (a): Increasing $k^{\mathrm{opt}}$ in PKPO training solves more problems during Gemma2 RL. (b): A higher $k^{\mathrm{opt}}$ makes the model learn to have higher entropy during RL. Thus, by optimizing for $\mathrm{pass@k}$ with $k>1$ instead of $\mathrm{pass@1}$, the model tends to have higher entropy leading to better exploration and solving more problems. Note that the size of one epoch, which is 750 steps, is evident in (a), where we see the slope decrease at each epoch boundary.
(a) $\mathrm{pass}@1$ ($k^{\mathrm{eval}}=1$)
(b) $\mathrm{pass}@2$ ($k^{\mathrm{eval}}=2$)
(c) $\mathrm{pass}@4$ ($k^{\mathrm{eval}}=4$)
(d) $\mathrm{pass}@8$ ($k^{\mathrm{eval}}=8$)
(e) $\mathrm{pass}@12$ ($k^{\mathrm{eval}}=12$)
(f) $\mathrm{pass}@16$ ($k^{\mathrm{eval}}=16$)
Figure 7: Effect of $k^{\mathrm{opt}}$ (used in our PKPO training) on the rolling $\mathrm{pass@k^{eval}}$ in Gemma2 RL. Setting $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ usually achieves the best $\mathrm{pass@k^{eval}}$. Prior work [39] (which is equivalent to the specific case of $k^{\mathrm{opt}}=n=16$ in our notation) is also shown for comparison, and suffers here presumably due to the larger estimator variance and unreliable gradient (see also Figure 4).
(a) Cumulative solve rate
(b) $\mathrm{pass}@1$ ($k^{\mathrm{eval}}=1$)
(c) $\mathrm{pass}@16$ ($k^{\mathrm{eval}}=16$)
Table 7: Results for Gemma2-2B on the MATH benchmark.
Gemma2-2B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 15.91 $\pm$ 0.40 18.12 $\pm$ 0.43 23.37 $\pm$ 0.48 29.58 $\pm$ 0.53 35.02 $\pm$ 0.58
k_opt=2 15.15 $\pm$ 0.41 20.73 $\pm$ 0.45 25.81 $\pm$ 0.50 31.96 $\pm$ 0.55 37.75 $\pm$ 0.60
k_opt=4 14.86 $\pm$ 0.43 19.86 $\pm$ 0.47 27.59 $\pm$ 0.52 34.27 $\pm$ 0.58 38.91 $\pm$ 0.62
k_opt=8 14.19 $\pm$ 0.46 19.33 $\pm$ 0.50 26.45 $\pm$ 0.55 35.49 $\pm$ 0.60 40.73 $\pm$ 0.65
[39] 13.11 $\pm$ 0.50 18.09 $\pm$ 0.54 24.58 $\pm$ 0.60 31.81 $\pm$ 0.66 37.24 $\pm$ 0.71
EntropyReg 14.51 $\pm$ 0.48 18.95 $\pm$ 0.52 25.33 $\pm$ 0.58 30.95 $\pm$ 0.64 36.18 $\pm$ 0.69
Table 8: Results for Gemma2-2B on the Coding benchmark.
Gemma2-2B k_eval=1 k_eval=2 k_eval=4 k_eval=8 k_eval=16
k_opt=1 19.82 $\pm$ 0.53 23.81 $\pm$ 0.57 29.75 $\pm$ 0.62 36.33 $\pm$ 0.66 42.04 $\pm$ 0.71
k_opt=2 18.70 $\pm$ 0.54 26.94 $\pm$ 0.59 33.82 $\pm$ 0.64 40.95 $\pm$ 0.69 47.03 $\pm$ 0.74
k_opt=4 18.69 $\pm$ 0.56 26.43 $\pm$ 0.61 36.81 $\pm$ 0.67 44.81 $\pm$ 0.73 50.55 $\pm$ 0.78
k_opt=8 17.94 $\pm$ 0.59 25.86 $\pm$ 0.64 35.88 $\pm$ 0.70 46.45 $\pm$ 0.77 52.83 $\pm$ 0.83
[39] 16.81 $\pm$ 0.65 24.27 $\pm$ 0.69 33.11 $\pm$ 0.76 41.98 $\pm$ 0.84 47.26 $\pm$ 0.89
EntropyReg 18.05 $\pm$ 0.62 25.13 $\pm$ 0.67 34.01 $\pm$ 0.74 40.88 $\pm$ 0.81 46.15 $\pm$ 0.86